What are the three longest wavelengths for standing sound waves in a 122-cm-long tube under the following conditions.

a)The tube is open at both ends.
b)The tube is open at one end, closed at the other?
Answer #1
given that
length of tube = L = 122cm = 1.22m
when the tube is open at both ends
first harmonic wave length = λ1= 2L = 2*1.22 = 2.44m
second = λ2= L = 1.22m
third = λ3= 2/3*L = 2/3 * 1.22 =0.81m
     2.44m,1.22m,0.81m
when the tube is open at one end
first harmonic = λ1= 4L = 4 * 1.22=4.88m
third harmonic = λ3= 4L/3= 4.88/3 =1.62m
fifth harmonic = λ5= 4L/5 = 0.976m
    4.88m,1.62m,0.976m